The sound wave has frequency 500Hz and velocity 350m/s. What is the distance between the two particles having phase difference of 60 degree? 
A. 0.7 cm
B. 12 cm
C. 120 cm
D. 70 cm 

Grade:12th pass

1 Answers

jayanti
4 years ago
Given , velocity = 350m/s
frequency = 500 Hz
so , wavelength = velocity / frequency = 350/500 = 0.7m
given phase difference = 60 = pie /3
phase difference = 2 pie x/ wavelength 
x = phase difference X wavelength / 2 pie
    pie 0.7/ 3 X 2 pie = 0.116m = 12 cm